OpenStudy (anonymous):

Write the expression sec (sin^-1 x) as an algebraic expression in x (without trig or inverse trig functions).

OpenStudy (dumbcow):

the key is to convert "sec" into expression with "sine" \[\sec x = \frac{1}{\cos x} = \frac{1}{\sqrt{1 - \sin^{2} x}}\] since \[\sin (\sin^{-1} x) = x\] \[\sec(\sin^{-1} ) = \frac{1}{\sqrt{1-x^2}}\]

OpenStudy (dumbcow):

another approach is to say \[x = \sin \theta\] then find sec(theta) using triangle 

OpenStudy (anonymous):

ahh makes sense. thank you. i have a similar problem sin(sin^-1 x + cos^-1 x) rewrite as an expression in x

OpenStudy (anonymous):

if you could help me on this that would be great

OpenStudy (dumbcow):

first use angle sum identity \[\sin(a +b) = \sin a \cos b + \sin b \cos a\] \[\sin(a) = \cos(b) = x\] \[\sin b = \sin(\cos^{-1} x) = \sqrt{1-x^2}\] \[\cos a= \cos(\sin^{-1} x) = \sqrt{1-x^2}\] *I hope the a,b notation didnt confuse you :)

OpenStudy (anonymous):

that helps. thanks